In a duopoly market, two firms produce the identical products, the cost function of firm 1 is: C1=20q1, the cost function of firm 2 is: C2=20q2 , the market demand function is: P=500-2Q , here Q = q1+q2


In a Bertrand model, the two firms set their price simultaneously, assume both firms do not have production capacity limits, and there is no collusion. What is the market equilibrium price and quantity?


If the two firms decide to form a Cartel, i.e. they want to maximize the profit of the whole industry, and then split the production and profit evenly. What is the market price? What is the industry’s total quantity produced? What is the quantity produced and profit of each firm?


1
Expert's answer
2021-12-28T08:42:06-0500

Solution:

The market equilibrium price and quantity in a Bertrand model:

TR1 = P "\times" Q1 = (500 – 2Q1 + Q2) "\times" Q1 = 500Q1 – 2Q12 + Q2Q1

MR1 = "\frac{\partial TR_{1} } {\partial Q_{1} }" = 500 – 4Q1 + Q2

Set MR1 to MC1:

500 – 4Q1 + Q2 = 20

500 – 20 - 4Q1 + Q2 = 0

480 - 4Q1 + Q2 = 0

Q1 = 120 + 0.25Q2

 

TR2 = P "\times" Q2 = (500 – 2Q1 + Q2) "\times" Q2 = 500Q2 – 2Q1Q2 + Q22

MR2 = "\frac{\partial TR_{2} } {\partial Q_{2} }" = 500 – 2Q1 + 2Q2

Set MR2 to MC2:

500 – 2Q1 + 2Q2 = 20

500 – 20 - 2Q1 + 2Q2 = 0

480 - 2Q1 + 2Q2 = 0

240 – Q1 + Q2 = 0

Q2 = 240 + Q1

 

Q1 = 120 + 0.25Q2

Q1 = 120 + 0.25(240 + Q1)

Q1 = 120 + 60 + 0.25Q1

Q1 – 0.25Q1 = 180

0.75Q1 = 180

Q1 = 240

 

Q2 = 240 + Q1 = 240 + 240 = 480

Market equilibrium quantity = 240 + 480 = 720

Substitute to derive the market equilibrium price:

P=500-2Q = 500 – 2(720) = 500 – 1440 = -940